Mary and Jane shop at an office supplies store. Mary spends $20 on 10 pens and
10 notebooks. Jane spends $15 on 5 pens and 8 notebooks. Find the prices of the
stationery.

Answer by mananth(16946)

let price of pen be $x
price of notebook be $y/10
...
10x+10y=20
x+y=2.......................1
5x+8y=15....................2
multiply equation 1 by -5 and add it to equation2
-5x-5y=-10
add to 2
8y-5y=15-10
3y=5
/3
y = 5/3
...
plug value of y in equation 1
x+5/3 = 2
(3x+5)/3=2
3x+5 =6
3x=6-5
3x=1
x= 0.33
...
Pen costs 33 cents
notebook costs $1.67
...
m.ananth@hotmail.ca